# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

### v1

@@ Abstract
The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and the underutilization of spectral complementarity. Existing methods often fail to decouple shared structural features from modality-specific radiometric attributes, leading to feature conflicts and information loss. To address this issue, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-specific) components in the Fourier domain. Specifically, PAD consists of two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features through convolution-guided scaling to enhance geometric consistency, and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency details and low-frequency structures using frequency-adaptive multilayer perceptrons. This approach leverages SAR’s sensitivity to morphological features and RGB’s spectral richness. Extensive experiments on WHU-OPT-SAR and DDHR-SK datasets demonstrate state-of-the-art performance. Our work establishes a new paradigm for physics-aware multi-modal fusion in remote sensing. The code will be available at https://github.com/RanFeng2/PAD.
@@ I Introduction
Land cover classification (LCC) is a fundamental task in Earth observation that involves categorizing various surface materials, such as farmlands, forests, water bodies, and urban areas, using remote sensing imagery. This capability is essential for environmental monitoring, resource management, and urban development [1]. In recent years, fusing Synthetic Aperture Radar (SAR) and RGB imagery has emerged as a promising approach to improve LCC performance. SAR provides robust structural information that is less susceptible to atmospheric and illumination variations, while RGB imagery offers rich spectral details, compensating for SAR-specific artifacts such as speckle noise [2, 3].
Despite their complementary relation, a key challenge remains: the inherent imaging heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB), which leads to significant spectral misalignment and hinders the effective fusion of cross-modal features. Compared to conventional multi-modal fusion in natural images, integrating SAR and RGB data presents more pronounced radiometric and geometric disparities, limiting the performance of existing frameworks [4, 5, 6].
@@
Figure 1: SAR-RGB Spatial-Frequency Complementarity Analysis. Frequency-domain representations capture modality complementarity more effectively than spatial features. (a) Spatial discrepancies exhibit anisotropic distributions; (b) relative amplitude differences are pronounced in high-frequency regions; (c) phase discrepancies are minimal and clustered in low-frequency bands; (e)–(f) variations in sampling rates significantly affect the spatial discrepancy distribution. Analysis conducted on WHU-OPT-SAR (see Section III-A).
Deep learning methods, particularly convolutional neural networks (CNNs), have achieved notable success in LCC tasks due to their powerful feature extraction capabilities [7]. To mitigate modality heterogeneity in SAR-RGB fusion, most approaches employ dual-branch architectures to independently extract multi-scale features from each modality [8], followed by fusion via simple operations such as element-wise summation, mean aggregation, or channel-wise concatenation, before processing with a shared decoder or segmentation head. Although this fusion scheme relies on accurately modeling complementary features, the intrinsic heterogeneity between SAR and RGB often impedes the extraction of truly complementary representations. Recent studies, inspired by contrastive learning [9, 10, 11], have explored aligning highly correlated, modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts. For example, Li et al. [12] considered phase as a modality-shared feature and used it as input in the intermediate network stage. However, they not only disregarded high-frequency amplitude patterns, which are essential for terrain delineation, thus limiting the effectiveness of the fusion, but also failed to systematically prove the complementarity of phase between the two modalities.
Through comprehensive spectral analysis (Fig. 1), we identify three critical frequency-domain priors that have been previously neglected: 1) Spatial-domain alignment (Fig. 1(a)) is significantly hampered by anisotropic discrepancies between modalities, rendering direct pixel- or feature-level matching unreliable; 2) Existing phase-based strategies overlook the inevitable slight differences between different frequency bands (Fig. 1(c)), leading to information loss; and 3) Amplitude mismatches between SAR and RGB modalities present different patterns in the high/low-frequency regions, most pronounced in the high-frequency band (Fig. 1(b)), highlighting the need for frequency-aware difference modeling.
@@
Motivated by these insights, we propose a novel fusion method, Phase-Amplitude Decoupling (PAD). The central idea behind PAD is to disentangle SAR and RGB modalities into phase components, which capture modality-invariant structural information, and amplitude components, which encode modality-specific details. By explicitly aligning the shared phase representations, PAD enhances the extraction of complementary amplitude features, thereby reducing feature conflicts caused by cross-modal heterogeneity. Specifically, PAD comprises two key modules: 1) a Phase Spectrum Correction (PSC) module that ensures robust phase alignment across modalities, and 2) an Amplitude Spectrum Fusion (ASF) module that selectively integrates complementary amplitude cues for improved fusion performance.
Our contributions are summarized fourfold:
@@
  • Through the systematic analysis of conventional registration datasets, we establish the frequency distribution law of SAR-RGB common patterns (phase) and unique patterns (amplitude) for the first time, providing an interpretable physical prior for multi-source remote sensing fusion.
  • We propose a novel spectral decomposition method that orthogonally separates phase (modality-shared) and amplitude (modality-specific) features, effectively resolving feature conflicts in SAR-RGB fusion while maintaining geometric integrity.
@@
  • To address phase inconsistencies, the PSC module employs convolution-guided multiplicative scaling to align and enhance critical structural details, ensuring robust fusion.
@@
  • The ASF module leverages frequency-domain multilayer perceptrons to dynamically integrate high-frequency (local details) and low-frequency (global structure) components, capitalizing on the complementary strengths of SAR and RGB data for improved segmentation.
@@
Figure 2: Flowchart of the PAD Framework. The framework comprises the (a) SCF module, (c) ASF module, and (d) PSC module. Registered RGB and SAR images, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are first input, and their features, $x_{RGB}$ and $x_{SAR}$, are extracted at each backbone stage before being fed into $n$ PAD fusion modules. Meanwhile, (b) FD and (e) FR denote the transformations from the spatial to the frequency domain and vice versa, respectively. Each stage’s final output is fed into a shared decoder. Here, $\mathcal{A}$ and $\mathcal{P}$ represent amplitude and phase, and $C$ (outer circle) indicates channel-wise concatenation.
@@ II-A Multi-modal Fusion for Land Cover Classification
While single-modality approaches [13, 14, 15, 16, 17, 18, 19, 20] have achieved reasonable performance in homogeneous data scenarios, their inability to leverage cross-modal synergies limits their discriminative power in complex environments. SAR-RGB fusion methods generally follow two paradigms: knowledge transfer via domain adaptation and hierarchical feature integration. In this section, we analyze their technical evolution and inherent limitations.
@@ II-A1 Cross-Modal Transfer Learning
Transfer learning [21] seeks to adapt knowledge from one modality to another. However, the significant differences between SAR and RGB data often degrade learned representations, limiting their applicability in multi-modal tasks.
@@ II-A2 Hierarchical Feature Fusion
Feature fusion strategies have progressed through three developmental phases: a) Basic Fusion Operators: Early methods utilized basic operations such as concatenation or weighted summation [22]. However, these approaches are prone to modality dominance and information redundancy in complex scenes. b) Attention-Guided Fusion: Spatial-channel attention mechanisms were introduced to address these issues [23, 24]. For instance, MCANet introduces cross-modal attention (MCAM) to model inter-modality dependencies, MRFS [25] develops progressive cycle attention (PC-Att) for infrared-visible fusion, and CMX [26] introduces bidirectional channel-spatial attention to refine cross-modal features. However, most existing attention paradigms have high computational costs. c) Complementarity-Aware Fusion: Recent state-of-the-art approaches explicitly model modality-specific and modality-shared components. ASANet [27] uses semantic focusing modules (SFM) to emphasize complementary features, CDDFuse [10] applies contrastive learning to disentangle private and shared representations, and CEN [28] proposes a method that achieves complementary perceptual fusion through dynamic channel exchange and parameter-sharing-independent batch normalization. Furthermore, recent work [12] pioneers phase-structure alignment by leveraging phase information as a cross-modal invariant feature. Despite these advancements, spatial-domain methods still struggle with geometric misalignment and computational inefficiency.
@@ II-B Frequency Representation Learning in Vision Tasks
Frequency analysis offers powerful tools to address challenges in spatial fusion [29, 30, 31]. Three primary application directions are discussed below:
II-B1 Phase as a Modality-Invariant Representation
Decoupling amplitude and phase spectra has proven effective in domain generalization (DG), where models trained on multiple source domains must generalize to unseen target domains. In DG, phase spectra encode structural semantics (e.g., object shapes and edges), while amplitude spectra capture modality-specific styles (e.g., illumination and texture patterns) [32, 33]. Early DG methods [34, 35] attempted to disentangle style and content by normalizing feature statistics (mean/variance), assuming that style corresponds to statistical variations. However, such heuristic separations can distort structural content. Recent approaches employ Fourier transforms as explicit style-content separators. For example, amplitude-phase recombination [36] swaps the amplitude spectrum of a source image with that of a target domain while preserving the source phase, thereby synthesizing cross-domain training samples. This strategy compels networks to focus on phase-centric structural features, improving robustness to style shifts. Building on this, decompose-adjust-compose (DAC) [37] integrates Phase-Amplitude Decoupling with feature normalization to dynamically balance style and content in the frequency domain. These methods highlight a critical advantage of spectral decomposition: phase alignment preserves structural consistency across modalities, while amplitude manipulation enables adaptive style fusion.
II-B2 Frequency-Driven Global Context Modeling
Beyond style-content separation, the Fourier transform’s global receptive field has been exploited to overcome two key limitations of spatial-domain networks: restricted long-range dependency modeling and the quadratic complexity of self-attention. For instance, fast Fourier convolution (FFC) [38] introduced hybrid spatial-frequency processing with parallel branches, capturing global patterns via spectral operations while preserving local details through convolution. These principles have also been applied to optimize attention mechanisms. FFTformer [39] replaces dot-product attention with spectral element-wise multiplication, reducing complexity from $O(N^{2})$ to $O(N\log N)$ while maintaining global interactions. Similarly, GFNet [40] and AFNO [41] employ learnable frequency filters as token mixers, achieving significant speedups compared to efficient self-attention variants. These methods demonstrate that global context aggregation in the frequency domain is not only computationally efficient but also capable of fusing multi-scale information.
II-B3 Frequency Priors as Inductive Biases
Recent studies have shown that frequency-domain representations can be explicitly designed as task-specific inductive biases [42, 43, 44]. For DG, DFF [45] exploits the assumption that low frequencies control cross-domain structural consistency, introducing a frequency-domain dynamic filtering mechanism to amplify domain-invariant components while suppressing high-frequency variations. In dense prediction tasks, FreqFusion [46] applies adaptive frequency filtering: an Adaptive Low-Pass Filter (ALPF) generator smooths high-frequency perturbations to homogenize intra-class features, while an Adaptive High-Pass Filter (AHPF) generator restores critical boundary information lost during downsampling. This strategy encodes spectral priors that low frequencies ensure semantic consistency, while high frequencies enhance boundary discriminability. In the context of RGB-NIR denoising, FCENet [47] utilizes a Frequency Dynamic Selection Module (FDSM) to select useful frequency components across modalities and a Frequency Exhaustive Fusion Module (FEFM) to refine shared and modality-specific spectral features. These approaches demonstrate that decoupling and resetting frequency components effectively encode task-related physical principles, such as structural invariance and boundary sensitivity, injecting strong inductive biases into deep networks.
@@ III Proposed Method
This paper introduces a multi-modal fusion framework based on Phase-Amplitude Decoupling (PAD) for land cover classification using SAR and RGB images. PAD effectively integrates complementary information from both modalities while mitigating heterogeneous feature conflicts that often arise with traditional spatial fusion techniques. The overall architecture, depicted in Fig. 2, consists of three main modules: the Amplitude Spectrum Fusion (ASF) module, the Phase Spectrum Correction (PSC) module, and the Spatial Complementary Fusion (SCF) module.
@@ 17 $\mathrm{APPD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{P}^{(i)}\}_{i=1}^{N})$
III-A Frequency-domain Analysis for Modality Decoupling
To validate the physical significance of frequency-domain decoupling in multi-modal data, we conducted a systematic spectral analysis on WHU-OPT-SAR. This analysis includes visualizations of the averaged difference spectra (Fig. 1(a)–(c)) computed using Algorithm 1 to compare the differences between two modalities in different domains. The difference spectra are defined as:
@@ III-A Frequency-domain Analysis for Modality Decoupling
where $I_{\mathrm{SAR}}$, $\mathcal{A}_{\mathrm{SAR}}$, and $\mathcal{P}_{\mathrm{SAR}}$ denote the pixel values, amplitude, and phase spectra of SAR images, respectively; $I_{\mathrm{RGB}}$, $\mathcal{A}_{\mathrm{RGB}}$, and $\mathcal{P}_{\mathrm{RGB}}$ correspond to those of the RGB images; $\arg(\cdot)$ computes the angle of a complex number; ||$\cdot$|| ⋅ | denotes the absolute value; and $\epsilon$ is a small constant to prevent division by zero. By leveraging the periodicity of complex exponentials, APPD maps phase differences into the principal value range $[-\pi,\pi)$, eliminating abnormally large differences caused by phase wrapping (e.g., a true phase difference of $2\pi$ is corrected to 00).
Figure 1 reveals three key observations: (1) Spatial differences (RSD) exhibit anisotropic distributions; (2) Amplitude differences (RAD) predominantly emerge in high-frequency regions; (3) Phase differences (APPD) remain minimal with slight low-frequency clustering. The frequency-domain metrics (RAD, APPD) demonstrate superior specificity compared to spatial-domain RSD, confirming that spectral decomposition better captures cross-modal complementarity.
These analyses lead to two fundamental insights that drive our architecture design: a) Phase Sharing: The consistently low APPD values indicate strong phase consistency between modalities. This suggests that phase spectra contain shared information that can guide cross-modal learning. Our PSC Module thus performs pixel-level phase alignment while preserving this inherent consistency. b) Amplitude Complementarity: The RAD distribution reveals frequency-dependent amplitude divergence, with high-frequency components showing greater discrepancy than low-frequency ones. This localized spectral pattern motivates our ASF Module, which adjusts the frequency radius dynamically to handle frequency-specific amplitude variations.
These phenomena can be physically interpreted based on the inherent characteristics of each imaging modality. Specifically, phase discrepancies manifest differently across frequency bands. Low-frequency disparities primarily stem from geometric distortions in imaging, leading to overall deviation between the two modalities that cannot be registered, whereas high-frequency variations arise from SAR speckle noise affecting fine details. On the other hand, amplitude disparities emerge from varying sensitivities to different ground objects, with low/high frequency contrasts reflecting modality-specific responses to large/small scale structures (e.g., farmlands and forests vs. roads).
We further validate our spectral analysis through sensitivity experiments on sampling rates (Fig. 1(e)–(f)). By analyzing the mean and variance of the difference spectra, we found that both RAD and APPD metrics are largely invariant to sampling rate, confirming that the observed phase-amplitude complementarity originates from intrinsic modality properties, rather than from sampling artifacts.
TABLE I: Statistical Analysis Results for the Average Periodic Phase Difference Spectrum.
Metric Low Freq High Freq Whole Image
Shapiro-Wilk (Stat) 1.0044 1.0050 1.0131
Shapiro-Wilk (p) 1.0 1.0 1.0
@@
Figure 3: Statistical Analysis of APPD. (a) APPD spectrum; (b)–(d) Frequency histograms with Gaussian fits for low-frequency (LF, radius = 0.5 $\times$ half-diagonal), high-frequency (HF), and full spectra (ALL); (e)–(f) Mean-variance trajectories across sampling rates. The HF, LF, and ALL spectra show near-Gaussian invariance, while LF/ALL deviates slightly due to the dominant structural noise in the LF range. These results suggest that phase, especially HF, can serve as a robust shared feature.
To assess the statistical consistency of cross-modal phase spectra, we conducted normality tests (Shapiro-Wilk), along with skewness and kurtosis analyses (Table I). Frequency histograms and mean-variance trajectories (Fig. 3) indicate that the frequency difference spectra follow a near-Gaussian distribution, particularly for high-frequency components, whereas low-frequency components exhibit minor systematic deviations, likely due to structural noise.
Figure 4: Visualization of Spatial and Frequency Spectrum Differences in a Scenario with Missing RGB Regions. (a)–(b) show RGB images with missing regions; (d)–(f) present the difference spectrum when the missing regions are included; and (g)–(i) display the difference spectrum after excluding the missing regions. (c) indicates that variation in RSD is significantly higher than in RAD and APPD, demonstrating that frequency-domain representations are more robust to missing data.
Figure 5: Visualization of Spatial and Frequency Spectrum Differences in a Cloud-Obscured Observations Scenario. (a)–(c) present an example of cloud-obscured observations, with cloud masks generated using RGB, SAR, and NIR images from WHU-OPT-SAR [48]. (d) compares the stability of RSD, RAD, and APPD between cloud-covered and cloud-free regions. Numbers next to the arrow indicates the value of decline.
We also evaluate phase spectrum invariance under challenging conditions through two controlled experiments: (1) partial RGB data missing (Fig. 4) and (2) cloud-obscured observations (Fig. 5). In the missing RGB scenario, RSD is highly sensitive to data incompleteness (Fig. 4(d) vs. (g)), while phase differences remain consistent (Fig. 4(f) vs. (i)). Similarly, in cloud-affected cases, the APPD variance ($\Delta\approx 10^{-0.037}\%$) is negligible compared to that of RSD ($\Delta\approx 10^{5.81}\%$) and RAD ($\Delta\approx 10^{0.58}\%$) as shown in Fig. 5(d). Collectively, these results demonstrate that phase information (particularly in the high-frequency region) exhibits robust consistency between SAR and RGB modalities, supporting its role as a shared feature to mitigate inter-modal conflicts and leverage complementary information in multi-modal frameworks.

### v2

@@ Abstract
The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and underutilized spectral complementarity. Existing methods often fail to decouple shared structural features from modality-complementary radiometric attributes, causing feature conflicts and information loss. To address this, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-complementary) components in the Fourier domain, thus reinforcing shared structures while preserving complementary characteristics to improve fusion quality. Unlike prior approaches that overlook the distinct physical properties encoded in frequency spectra, PAD is the first to introduce explicit amplitude-phase decoupling for multi-modal fusion. Specifically, PAD comprises two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features via convolution-guided scaling to enhance geometric consistency; and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive multilayer perceptrons, leveraging SAR’s morphological sensitivity and RGB’s spectral richness. Extensive experiments on WHU-OPT-SAR and DDHR-SK datasets demonstrate state-of-the-art performance. Our work establishes a new paradigm for physics-aware multi-modal fusion in remote sensing. The code will be available at https://github.com/RanFeng2/PAD.
@@ I Introduction
Land cover classification (LCC) is a fundamental Earth observation task that categorizes surface materials, such as farmlands, forests, water bodies, and urban areas, using remote sensing imagery. This capability is essential for environmental monitoring, resource management, and urban development [1]. Recently, fusing Synthetic Aperture Radar (SAR) and RGB imagery has emerged as a promising approach to enhance LCC performance: SAR provides robust structural information resilient to atmospheric and illumination variations, while RGB imagery offers rich spectral details that mitigate SAR artifacts such as speckle noise [2, 3].
@@
Figure 1: Motivation and SAR-RGB Spatial-Frequency Difference Analysis. (a) SAR and RGB imagery exhibit modality gaps due to distinct sensing mechanisms (backscattering vs. reflectance). (b) Fusion feature space comprises heterogeneous, shared, and complementary components; effective fusion suppresses heterogeneous features while enhancing shared and complementary patterns. (c)-(g) Frequency-domain analysis on WHU-OPT-SAR (see Section III-A): (c) spatial discrepancies exhibit anisotropic distributions; (d) relative amplitude differences concentrate in high-frequency regions; (e) phase discrepancies are minimal and cluster in low-frequency bands; (f)-(g) sampling-rate variations significantly affect spatial discrepancy distributions.
Despite this complementarity, RGB-SAR fusion faces a key challenge: inherent imaging heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB) leads to significant spectral misalignment and hinders effective cross-modal feature integration (see Fig. 1(a)). In SAR images, objects often appear as discrete backscatter responses, lacking fundamental visual cues, such as geometric contours, textures, and consistent shapes, that are readily available in RGB imagery [4]. Compared to natural-image fusion, integrating SAR and RGB data presents pronounced radiometric and geometric disparities, limiting existing frameworks [5, 6, 7].
Deep learning, particularly convolutional neural networks (CNNs), has achieved notable success in LCC by extracting powerful features [8, 9]. To mitigate modality heterogeneity, most approaches use dual-branch architectures to extract multi-scale features independently, then fuse them via element-wise summation, mean aggregation, or channel-wise concatenation before a shared decoder or segmentation head [10]. However, accurate modeling of complementary features is still impeded by SAR-RGB heterogeneity. Inspired by contrastive learning [11, 12, 13], recent works align modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts (see Fig. 1(b)), yet spatial-domain modeling remains inefficient.
To better capture RGB-SAR complementarity, some methods explore frequency-domain fusion [14, 15, 16]. For example, Li et al. [17] treat phase as modality-shared features but discard amplitude information essential for terrain delineation and do not systematically validate phase sharing.
We conduct comprehensive spectral analysis (see Fig. 1(c)-(g)) to reveal three critical frequency-domain priors: 1) SAR-RGB spatial-domain differences are anisotropic and susceptible to pixel-level mismatch (see Fig. 1(c)); 2) average phase spectra differ minimally and mainly in low frequencies, indicating phase as shared information (see Fig. 1(e)); and 3) amplitude mismatches exhibit distinct high/low-frequency patterns, most pronounced in the high-frequency band, highlighting the need for frequency-aware modeling (see Fig. 1(d)).
@@
Motivated by these insights, we propose Phase-Amplitude Decoupling (PAD), which disentangles phase components capturing modality-invariant structures from amplitude components encoding modality-specific details. By explicitly aligning shared phase representations, PAD enhances complementary amplitude feature extraction and reduces cross-modal conflicts. PAD comprises two modules: Phase Spectrum Correction (PSC), which aligns and reinforces structural details via convolution-guided scaling; and Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive MLPs.
Our contributions are fourfold:
@@
  • We systematically analyze registration datasets to establish frequency distributions of SAR-RGB common (phase) and unique (amplitude) patterns, offering interpretable physical priors for multi-source fusion.
  • We introduce frequency-domain amplitude-phase decoupling for multi-modal fusion, orthogonally separating phase (modality-shared) and amplitude (modality-specific) features to resolve heterogeneity-induced conflicts.
@@
  • PSC employs convolution-guided multiplicative scaling to align and reinforce critical structural details for robust fusion.
@@
  • ASF uses frequency-domain MLPs to integrate complementary high-frequency (local details) and low-frequency (global structure) components, leveraging SAR’s morphological sensitivity and RGB’s spectral richness for improved segmentation.
@@
Figure 2: Flowchart of the PAD Framework. Registered RGB and SAR images, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are processed by asymmetric backbones to extract features $x_{\mathrm{RGB}}$ and $x_{\mathrm{SAR}}$ at each stage. These features are passed through $n$ PAD fusion modules, each comprising: (a) the SCF module, (c) the ASF module, and (d) the PSC module. (b) FD and (e) FR denote the spatial-to-frequency and frequency-to-spatial transformations, respectively. Each stage’s fused features are concatenated channel-wise (outer circle $C$) and fed into a shared decoder. Here, $\mathcal{A}$ and $\mathcal{P}$ represent amplitude and phase spectra.
@@ II-A Multi-modal Fusion for Land Cover Classification
Single-modality approaches [18, 19, 20, 21, 22, 23, 24, 25] perform well on homogeneous data but cannot exploit cross-modal synergies in complex environments. SAR-RGB fusion methods generally follow two paradigms: cross-modal transfer learning and hierarchical feature integration.
@@ II-A1 Cross-Modal Transfer Learning
Transfer learning [26] adapts knowledge from one modality to another, but the pronounced differences between SAR and RGB data often degrade learned representations, limiting multi-modal applicability.
@@ II-A2 Hierarchical Feature Fusion
Feature fusion has evolved through three stages: a) Basic Fusion Operators: Early methods used concatenation or weighted summation [27], but these are prone to modality dominance and redundancy. b) Attention-Guided Fusion: Spatial-channel attention mechanisms address these issues [28, 29]. For example, MCANet’s cross-modal attention (MCAM) models inter-modality dependencies; MRFS [30] introduces progressive cycle attention for infrared-visible fusion; and CMX [31] employs bidirectional channel-spatial attention. However, these paradigms incur high computational costs. c) Complementarity-Aware Fusion: Recent approaches explicitly disentangle shared and modality-specific features. ASANet [32] uses semantic focusing modules to emphasize complementary cues; CDDFuse [12] applies contrastive learning to separate private and shared representations; and CEN [33] leverages dynamic channel exchange with independent batch normalization. Despite these advances, spatial-domain methods still struggle with geometric misalignment and inefficiency.
@@ II-B Frequency Representation Learning in Vision Tasks
Frequency representation learning has emerged as a powerful tool across various vision tasks due to its capacity for global modeling and structural decomposition. In this subsection, we review three relevant areas:
II-B1 Spectral Decomposition for Structure-Style Separation
Fourier-based decomposition isolates phase (structural semantics) and amplitude (appearance) for domain generalization and style transfer [34, 35]. Early methods used feature statistics for style normalization [36, 37], while recent works employ Fourier transforms for clearer separation, enabling content-preserving augmentation by swapping amplitudes across domains [38, 39].
II-B2 Frequency-Driven Global Modeling
Fourier transforms provide a global receptive field useful for efficient attention and context modeling. Fast Fourier Convolution (FFC) [40] and FFTformer [41] leverage spectral processing to replace or accelerate spatial attention. Spectral token mixing has also been effective in lightweight transformers [42, 43].
II-B3 Frequency-Based Inductive Biases in Dense Prediction
Spectral priors enhance dense prediction tasks. DFF [44] uses frequency-domain dynamic filtering to amplify invariant low-frequency features, improving cross-domain generalization. FreqFusion [45] designs learnable high- or low-pass filters for semantic consistency and boundary refinement. FCENet [14] employs frequency-selective modules for cross-modal spectral correlations in RGB-NIR denoising.
In summary, frequency-based representations offer a flexible, physically interpretable design space for vision tasks; yet, few frameworks target task-specific fusion in remote sensing. We address this gap in the next section.
II-C Frequency-Domain Fusion Methods
Applying frequency representations to multi-modal fusion is recent. Shi et al. [15] introduce Frefusion, using spectral attention to merge modalities in the frequency domain. Frequency spectra naturally decompose signals into phase (modality-invariant structures) and amplitude (modality-specific details), offering a principled fusion framework.
However, existing methods often transplant spatial operators, MLPs [16], convolutions [14], or attention [15], into the spectral domain, modulating magnitude while neglecting the distinct roles of phase and amplitude. Consequently, shared information may be preserved, but complementary cues are suppressed.
In contrast, our PAD framework explicitly decouples phase and amplitude spectra. Through systematic spectral analysis (see Section III-A), we show that SAR and RGB share strong phase consistency while amplitude captures complementary patterns. PAD introduces (1) a Phase Spectrum Correction module for robust structural alignment and (2) an Amplitude Spectrum Fusion module with frequency-adaptive refinement. This decomposition-correction-fusion pipeline distinguishes PAD from previous FFT-based approaches by providing a physically interpretable strategy for multi-modal fusion.
The two most related works, FeINFN [16] and SFINet++ [46], also leverage amplitude and phase in the frequency domain for high-frequency detail enhancement and global context modeling. However, they lack a physically grounded interpretation of amplitude-phase roles and do not explicitly decouple shared versus modality-specific information. In contrast, PAD is the first framework to formulate and exploit amplitude-phase decomposition for principled alignment and complementary integration across modalities.
@@ III Proposed Method
To address the fundamental conflict between modality-invariant geometry and modality-specific appearance in SAR-RGB fusion, we introduce Phase-Amplitude Decoupling (PAD), a frequency-aware multi-modal fusion framework. The overall architecture, shown in Fig. 2, comprises three modules: the Spatial Complementary Fusion (SCF) module, which preserves fine-grained spatial priors; the Phase Spectrum Correction (PSC) module, which aligns and reinforces shared structural details; and the Amplitude Spectrum Fusion (ASF) module, which adaptively integrates complementary high- and low-frequency amplitude cues. Unlike previous methods that apply spectral operations uniformly, PAD explicitly decouples and recombines phase and amplitude with dedicated modules to enable physics-consistent fusion.
@@ 17 $\mathrm{APPD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{P}^{(i)}\}_{i=1}^{N})$
III-A Frequency-domain Priors for Amplitude-Phase Decoupling
To validate the physical significance of phase-amplitude decoupling and reveal the shared and complementary relations between RGB and SAR, we conducted a systematic spectral analysis on WHU-OPT-SAR dataset, the largest RGB-SAR segmentation benchmark.
Firstly, we visualized the averaged difference spectra (see Fig. 1(c)-(e)) using Algorithm 1 to compare modality differences in spatial, amplitude, and phase domains. This analysis shows that frequency-domain representations capture cross-modal sharing and complementarity more effectively than spatial representations.
The difference spectra are defined as:
@@ III-A Frequency-domain Priors for Amplitude-Phase Decoupling
where $I$, $\mathcal{A}$, and $\mathcal{P}$ denote pixel values, amplitude spectra, and phase spectra for SAR and RGB images; $\arg(\cdot)$ computes the complex angle; and $\epsilon$ prevents division by zero. By leveraging the periodicity of complex exponentials, APPD maps phase differences into $[-\pi,\pi)$, correcting wrap-around artifacts (e.g., $2\pi\to 0$).
Figure 1(c)-(e) reveals three priors: 1) Spatial differences (RSD) exhibit anisotropic distributions. 2) Amplitude differences (RAD) concentrate in high-frequency regions. 3) Phase differences (APPD) remain minimal, with slight low-frequency clustering. These frequency-domain metrics demonstrate superior explainability over spatial RSD, confirming that spectral decomposition better captures cross-modal specificity.
Physically, low-frequency phase disparities stem from geometric distortions (e.g., side-looking SAR vs. orthographic RGB), while high-frequency variations arise from SAR speckle noise. Amplitude disparities reflect modality-specific sensitivity to object scales: low frequencies encode large structures (farmlands, forests) and high frequencies capture fine details (roads).
These analyses yield two key insights that drive our architecture design: 1) Phase Sharing: Consistently low APPD values indicate strong phase consistency. The PSC module thus performs pixel-level alignment while preserving this shared information. 2) Amplitude Complementarity: RAD’s frequency-dependent divergence motivates ASF’s dynamic frequency-radius adjustment to handle amplitude variations.
Secondly, we conducted sensitivity experiments on sampling rates (see Fig. 1(f)-(g)) to evaluate the stability of our spectral metrics. By downsampling input images at various ratios and computing the mean and variance of RSD, RAD, and APPD across the dataset, we observed that RAD and APPD remain nearly constant, whereas RSD exhibits larger fluctuations. This invariance of RAD and APPD to sampling artifacts confirms that frequency-domain complementarity arises from intrinsic modality characteristics rather than from individual changes.
TABLE I: Statistical Analysis of the Average Periodic Phase Difference Spectrum.
Metric Low Frequency High Frequency Whole Image
Shapiro-Wilk Statistic 1.0044 1.0050 1.0131
Shapiro-Wilk p-value 1.0000 1.0000 1.0000
@@
Figure 3: Statistical Analysis of APPD. (a) APPD spectrum. (b)-(d) Frequency histograms with Gaussian fits for low-frequency (LF; radius = 0.5 $\times$ half-diagonal), high-frequency (HF), and all spectra (ALL). (e)-(f) Mean-variance trajectories across sampling rates. All spectra exhibit near-Gaussian invariance, with LF and ALL showing slight deviations due to structural noise in the low-frequency range, indicating that phase, particularly HF, can serve as a robust shared feature.
Thirdly, to assess the statistical consistency of cross-modal phase spectra, we performed normality tests (Shapiro-Wilk) and calculated skewness and kurtosis for APPD values at each frequency band (see Table I). Frequency histograms and mean-variance trajectories (see Fig. 3) demonstrate that APPD distributions closely follow a Gaussian profile, especially in high-frequency regions. Low-frequency components show minor deviations, likely due to structural noise, but overall the near-Gaussian behavior validates phase spectra as a reliable shared feature. In a word, phase is so similar between RGB and SAR that it is almost unaffected by individual differences.
Figure 4: Visualization of Spatial and Frequency Spectrum Differences under Missing RGB Regions. (a)-(b) RGB images with missing regions; (c) variation in RSD, RAD, and APPD, showing RSD’s greater sensitivity to missing data; (d)-(f) difference spectra including missing regions; (g)-(i) difference spectra excluding missing regions. These results demonstrate that frequency-domain metrics (RAD, APPD) are more robust to missing data than spatial-domain RSD.
Figure 5: Visualization of Spatial and Frequency Spectrum Differences under Cloud Obscuration. (a)-(c) Example observations with cloud masks derived from RGB, SAR, and NIR channels in WHU-OPT-SAR [47]. (d) Comparison of RSD, RAD, and APPD stability between cloud-covered and cloud-free areas; arrows denote the value of decline.
Finally, we evaluated phase invariance under two challenging scenarios: 1) partial RGB data removal (see Fig. 4) and 2) cloud-obscured observations (see Fig. 5). In the former, RSD increased dramatically when RGB pixels were missing (compare Fig. 4(d) and (g)), while APPD remained stable (compare Fig. 4(f) and (i)). In the latter, APPD variance ($\Delta\approx 10^{-0.037}\%$) was negligible compared to that of RSD ($\Delta\approx 10^{5.81}\%$) and RAD ($\Delta\approx 10^{0.58}\%$) as shown in Fig. 5(d). These results confirm that phase information, particularly at high frequencies, retains robust consistency across modalities, underscoring its role as a shared feature to mitigate inter-modal conflicts and enhance complementary fusion.